OpenStudy (anonymous):

It takes Trevon ten hours to clean an attic. Cody can clean the same attic in seven hours. Find how long it would take them if they worked together

OpenStudy (ybarrap):

Rate that Trevon can clean an attic = T= 1 attic/10hrs Rate that Cody can clean an attic = C= 1 attic/7hrs t = time it takes (in hours) for both of them together to clean 1 attic 1 = T*t + C+t, because in t hours, both of them working together will clean 1 attic solve for t -- that's it.

hero (hero):

T = hours it takes Trevon to clean attic = 10 C = hours it takes Cody to clean attic = 7 Then the time it takes both to clean the attic while working together is represented by: \[\frac{CT}{C + T} = t\] Solving for t... \[\frac{(7)(10)}{7 + 10} = \frac{70}{7} = 4\frac{2}{17}\text{hours}\]